OpenStudy (anonymous):

Solve the following logarithm equation. log (x + 2) – log(x – 5) = log 4

OpenStudy (anonymous):

is this log base ten?

OpenStudy (anonymous):

ooops no it is a minus sign, step one will be \[\log(a)-\log(b)=\log(\frac{a}{b})\]

OpenStudy (anonymous):

wrong again it is \[\log(4)\] on the other side so you have \[\log(\frac{x+2}{x-5})=\log(4)\] meaning \[\frac{x+2}{x-5}=4\]

OpenStudy (anonymous):

then solve via \[x+2=4(x-5)\] \[x+2=4x-20\] etc